Circular No. 6381

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions)

BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science) Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1996X IN NGC 5061 N. B. Suntzeff reports that spectrophotometry (range 370–730 nm,

resolution 0.35 nm) obtained by M.-T. Ruiz with the Cerro Tololo 4-m telescope on Apr. 14.2 UT confirms that this is a type-Ia supernova near maximum (cf. IAUC 6380). The spectrum is similar to that of SN 1981B at maximum, which is considered a typical type-Ia supernova. The spectrum shows some unresolved interstellar Na I D (0.03 nm total) and Ca II H (0.04 nm) and K (0.02 nm) absorption. The Galatic reddening to this galaxy is E(B−V) = 0.25 (Burstein and Heiles 1984, Ap.J. Suppl. 54, 33). D. Depoy estimates V = 13.3 on Apr. 14.1 UT with the Cerro Tololo 0.9-m telescope.

L. Wang, and J. C. Wheeler, University of Texas at Austin and Mc-Donald Observatory, report: “SN 1996X was observed on Apr. 14.3 UT using the 2.1-m telescope (+ Imaging Grism Instrument) at McDonald Observatory. Preliminary inspection of a spectrum (range 424.7–813.5 nm, resolution 2.0 nm) shows that SN 1996X is a type-Ia supernova before optical maximum.”

S. Benetti and F. Patat, European Southern Observatory (ESO); and J. Beuing report that inspection of a fully-reduced CCD spectrogram (range 465–667 nm, resolution 0.29 nm) obtained on Apr. 14.02 UT with the ESO 1.5-m telescope (+ Boller & Chivens spectrograph) confirms that SN 1996X is a type-Ia supernova, ∼ 1 week before maximum light. The spectrum is dominated by P-Cyg lines of intermediate-mass elements superimposed on a blue continuum. Visible lines include Si II at 505.1, 595.8, and (very strong) 635.5 nm; Si III at 574.0 nm; Fe III at 512.9 nm; and S II at 532.2, 546.8, 561.2, and 565.4 nm. The expansion velocity deduced from the minimum of Si II at 635.5 nm, corrected by the redshift (2032 km/s) of the parent galaxy, is ∼ 11 150 km/s. Preliminary photometry from CCD frames taken with the Dutch 0.9-m telescope at La Silla (ESO) on Apr. 14.29 gives V = 13.38, B − V = −0.02, U − B = −0.59, V − R = +0.09, V − I = +0.02.

Also, L. Germany and B. Schmidt, Mount Stromlo and Siding Spring Observatories (MSSSO), report that examination of a spectrum (range 330– 1000 nm) taken on Apr. 14.4 UT by K. Freeman (MSSSO) on the MSSSO 2.3-m telescope (+ Imager + Grism) shows this to be a type-Ia supernova ∼ 2 days past maximum. Likewise, P. Garnavich and R. Kirshner, Harvard-Smithsonian Center for Astrophysics, report that a spectrogram obtained by E. Barton with the 1.5-m Tillinghast telescope on Apr. 14.3 shows this to be a type-Ia supernova near maximum brightness.

1996 April 15  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1996 CBAT  Daniel W. E. Green